5.18 Assume a Poisson distribution.

a. Ifl = 2.5,findP1X = 22.
b. Ifl = 8.0,findP1X = 82.
c. Ifl = 0.5,findP1X = 12.

d. Ifl = 3.7,findP1X = 02.
Answer #1

Poisson distribution P(x-x) = ed a) If x=25 P(x - 2) = 6-205(2-5) 2! b) = 0.2565 If i=8.0 P(x=8) - 1-889 = 0.1395 c) 8-0.5 p(